# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

### v5

@@ Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability. The resultant differences are around $30\ \%$ for Ar and $80\ \%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated from the $3p$ shell electrons for Ar. For Xe, $5p$ states dominate the contribution, but, $4d$ states give $\sim 40\ \%$ contribution at 1 keV.
@@ 1 Introduction
The Migdal effect migdal1939ionizatsiya ; migdal1941ionizacia ; Landau:1991wop is effective for searching small energy signals from dark matter. The Migdal effect is a inelastic process of dark matter-nucleus scattering and causes the excitation/ionization of electrons from the nuclear recoil. The excitation/ionization of electrons returns electromagnetic signals that are easier to detect than heat in the case of elastic scattering. The Migdal effect is not observed in nuclear scattering, but, observed in $\alpha$ decay Rapaport:1975zza ; Rapaport:1975zz ; fischbeck1975spectroscopy ; fischbeck1977angular and $\beta$ decay boehm1954internal ; berlovich1965investigation ; couratin2012first ; Lienard:2015joa ; Fabian:2018pin . Now, experiments for detecting the Migdal effect from nuclear recoils are proceeding Nakamura:2020kex ; Araujo:2022wjh . Recently, the Migdal effect is studied or used in many papers baur1983ionisation ; vegh1983multiple ; wauters1997recoil ; berakdar2001electronic ; Vergados:2005dpd ; Moustakidis:2005gx ; Ejiri:2005aj ; talman2006excitations ; Bernabei:2007jz ; liertzer2012multielectron ; Vergados:2013raa ; pindzola2014neutron ; Sharma:2017fmo ; Ibe:2017yqa ; Dolan:2017xbu ; LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr ; Coskuner:2018are ; EDELWEISS:2019vjv ; Tvrznikova:2019tgx ; CDEX:2019hzn ; XENON:2019zpr ; Baxter:2019pnz ; Essig:2019xkx ; Liang:2019nnx ; Kozaczuk:2020uzb ; SENSEI:2020dpa ; Liu:2020pat ; Nakamura:2020kex ; Pindzola:2020whx ; Kahn:2020fef ; Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu ; LUX:2020yym ; Liang:2020ryg ; He:2020sat ; Flambaum:2020xxo ; Akerib:2021pfd ; Collar:2021fcl ; Bell:2021zkr ; Liao:2021yog ; Menshikov:2021jxa ; Acevedo:2021kly ; COSINE-100:2021poy ; Berlin:2021zbv ; CDEX:2021cll ; Wang:2021oha ; Bell:2021ihi ; Abbamonte:2022rfh ; Chatterjee:2022gbo ; EDELWEISS:2022ktt ; Angevaare:2022zhy ; Campbell-Deem:2022fqm ; Araujo:2022wjh ; DarkSide:2022dhx ; Blanco:2022pkt ; Cox:2022ekg ; GlobalArgonDarkMatter:2022ppc ; Tomar:2022ofh ; Adams:2022zvg ; Berghaus:2022pbu ; Li:2022acp ; NEWS-G:2023qwh ; DarkSide-50:2023fcw ; SuperCDMS:2023sql ; Bell:2023uvf ; Tilly:2023fhw ; AtzoriCorona:2023ais ; Qiao:2023pbw ; Xu:2023wev ; LZ:2023poo ; PandaX:2023xgl ; Yun:2023huf ; Gu:2023pfg ; Li:2023xkf ; Herrera:2023xun ; Kim:2023lkb ; Dzuba:2023zcq ; SENSEI:2023zdf ; Liang:2024tef ; He:2024hkr ; Balan:2024cmq ; MIGDAL:2024alc ; Kang:2024kec ; Ibarra:2024mpq ; Lee:2024wzd , which includes that the comparison LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr to bremsstrahlung of photon Kouvaris:2016afs ; Millar:2018hkv , theoretical comparison to the dark matter-electron scattering Baxter:2019pnz ; Essig:2019xkx , the Migdal-photoabsorption relation Liu:2020pat , formulation for semiconductor response Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu , and the semi-inclusive calculation Cox:2022ekg .
@@ 2 Energy eigenstates of atomic system
where $\vec{p}_{n}$ and $\vec{p}_{e,i}$ is the three-dimensional momentum of nucleus and $i$-th electron, $\vec{r}$ is the three-dimensional coordinates, $m_{n}$ and $m_{e}$ are the mass of nucleus and electron, $Z$ is the atomic number, and $\alpha$ is the fine-structure constant.
@@
The above Hamiltonian can be separated by the center of mass energy and the others. Therefore, the energy eigenstates of atom are given by $|\Psi_{A}\rangle=|\vec{p}_{\rm CM}\rangle\otimes|\Psi_{\rm re}\rangle$. Here, $|\Psi_{\rm re}\rangle$ is the energy eigenstate of the Hamiltonian for the relative coordinates
@@
and $|\vec{p}_{\rm CM}\rangle$ is the plane wave eigenstate for the center of mass state. Note that $H_{\rm re}$ includes the potential of nucleus being at the origin by the approximation. Therefore, the relative coordinates stand for the positions of electrons.
In the following, the energy eigenstate $|\Psi_{\rm re}\rangle$ is assumed to be given by the Slater determinant of single-electron wave function. Also, the single-electron wave functions are expressed by the spherical spinors (or spinor spherical harmonics) (see, e.g., Ref. johnson2007atomic for review).
@@ 3 Interaction potential
The interaction with the nucleus can be ignored approximately when the excitation and ionization processes are considered. This is because the potential from the nucleus interactions depend on the coordinates as $\vec{r}-\vec{r}_{n}$ and $\vec{r}_{n}$ is related to the relative coordinates $\vec{r}_{i}$ with the suppression factor of $\approx m_{e}/m_{n}$.
@@ 4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in},s_{\rm in}\rangle_{I/F}$. The incoming electron states can be taken as plane waves because the atom is regarded to be neutral far from the atom. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
@@
×$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I},s_{{\rm in},I}\rangle$
×$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F},s_{{\rm in},F}|\vec{r}\rangle e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}).$ (7)
@@
×$\times\left(\frac{m_{A}}{m_{e}}\right)^{2}(2\pi)^{4}\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$
×$\times\left|\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right|^{2}.$ (12)
@@
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the difference of ionization cross section is around $30\ \%$ from the database value in 5 keV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around $60\ \%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the difference of ionization cross section is around 80 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This gap dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term is effective at low energy and negligible at high energy. Note that the probability is almost dominated by the excitation/ionization from the outermost shell. For Ar, $3p$ states dominate the contribution and $2p$ and $3s$ states give $\sim 1\ \%$ contribution at $E_{{\rm kin},I}=10$ keV. For Xe, $5p$ states dominate up to $\sim$ 100 eV, but, $4d$ states have $\sim 40\ \%$ contribution at $E_{{\rm kin},I}=1$ keV.
@@ 5 Comparison with other results
where $\beta$ is the velocity of incoming electron and the values of $M_{\rm Salvat}^{2}$ and $C_{\rm Salvat}$ are taken from the figure in Ref. salvat2022inelastic . This calculation used the shell corrections, self-consistent Dirac-Hartree-Fock-Slater potential, independent-electron approximation, and relativistic plane-wave Born approximation. They reported less than 1 % precision for $E\gtrsim 100$ MeV and $\sim 8$ precision for $E\sim 5$ MeV for high $Z$ atoms and good agreement with the results of Ref. inokuti1975total ; inokuti1981oscillator .
@@
The binding energies and wave functions of electrons are calculated with electron correlations, e.g., Ref. Baxter:2019pnz shows $\sim 5$ precision for binding energies which is better than Ref. Ibe:2017yqa . Also, Ref. Cox:2022ekg shows $\sim 20$ differences of ionization probabilities for xenon 5s and argon 3p states with a 20 keV ionized electron. These uncertainties are significant for outer shell states and this comparison method actually relies on the precision of outermost shell states. Note that the uncertainty becomes much smaller for inner shell electrons. Density functional theory may provide better precision of binding energies, but, it does not provide physical wave functions.
@@ 6 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The differences are $\sim 30\ \%$ for Ar at 10 keV and $\sim 80\ \%$ for Xe compared with the database value shown in Fig. 1. For Xe, the database value is limited and the differences may be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. Also, the approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected differences for excitation are not reliable. The differences of the low energy threshold come from the kinematical constraint, $E_{{\rm kin},I}>2\Delta E$. Also, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ differences with experimental data Ibe:2017yqa .

### v6

@@ Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability. The resultant differences are around 30%30\ $\%$ for Ar and 80%80\ $\%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated from the $3p$ shell electrons for Ar. For Xe, $5p$ states dominate the contribution, but, $4d$ states give $\sim 40\ \%$ contribution at 1 keV.
@@ 1 Introduction
The Migdal effect migdal1939ionizatsiya ; migdal1941ionizacia ; Landau:1991wop is an inelastic process of scattered nucleus. It causes the excitation/ionization of electrons from the nuclear recoil or decay. The Migdal effect is observed in $\alpha$ decay Rapaport:1975zza ; Rapaport:1975zz ; fischbeck1975spectroscopy ; fischbeck1977angular and $\beta$ decay boehm1954internal ; berlovich1965investigation ; couratin2012first ; Lienard:2015joa ; Fabian:2018pin , however, not observed in nuclear scattering. Now, experiments for detecting the Migdal effect from gas probes are proceeding Nakamura:2020kex ; Araujo:2022wjh .
@@
The Migdal effect is also effective for searching small energy signals from dark matter. The Migdal effect occurs as an inelastic process of dark matter-nucleus scattering and causes the excitation/ionization of electrons from the nuclear recoil. The excitation/ionization of electrons returns electromagnetic signals that are easier to detect than heat in the case of elastic scattering. Recently, the Migdal effect is studied or used in many papers baur1983ionisation ; vegh1983multiple ; wauters1997recoil ; berakdar2001electronic ; Vergados:2005dpd ; Moustakidis:2005gx ; Ejiri:2005aj ; talman2006excitations ; Bernabei:2007jz ; liertzer2012multielectron ; Vergados:2013raa ; pindzola2014neutron ; Sharma:2017fmo ; Ibe:2017yqa ; Dolan:2017xbu ; LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr ; Coskuner:2018are ; EDELWEISS:2019vjv ; Tvrznikova:2019tgx ; CDEX:2019hzn ; XENON:2019zpr ; Baxter:2019pnz ; Essig:2019xkx ; Liang:2019nnx ; Kozaczuk:2020uzb ; SENSEI:2020dpa ; Liu:2020pat ; Nakamura:2020kex ; Pindzola:2020whx ; Kahn:2020fef ; Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu ; LUX:2020yym ; Liang:2020ryg ; He:2020sat ; Flambaum:2020xxo ; Akerib:2021pfd ; Collar:2021fcl ; Bell:2021zkr ; Liao:2021yog ; Menshikov:2021jxa ; Acevedo:2021kly ; COSINE-100:2021poy ; Berlin:2021zbv ; CDEX:2021cll ; Wang:2021oha ; Bell:2021ihi ; Abbamonte:2022rfh ; Chatterjee:2022gbo ; EDELWEISS:2022ktt ; Angevaare:2022zhy ; Campbell-Deem:2022fqm ; Araujo:2022wjh ; DarkSide:2022dhx ; Blanco:2022pkt ; Cox:2022ekg ; GlobalArgonDarkMatter:2022ppc ; Tomar:2022ofh ; Adams:2022zvg ; Berghaus:2022pbu ; Li:2022acp ; NEWS-G:2023qwh ; DarkSide-50:2023fcw ; SuperCDMS:2023sql ; Bell:2023uvf ; Tilly:2023fhw ; AtzoriCorona:2023ais ; Qiao:2023pbw ; Xu:2023wev ; LZ:2023poo ; PandaX:2023xgl ; Yun:2023huf ; Gu:2023pfg ; Li:2023xkf ; Herrera:2023xun ; Kim:2023lkb ; Dzuba:2023zcq ; SENSEI:2023zdf ; Liang:2024tef ; He:2024hkr ; Balan:2024cmq ; MIGDAL:2024alc ; Kang:2024kec ; Ibarra:2024mpq ; Lee:2024wzd , which includes that the comparison LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr to bremsstrahlung of photon Kouvaris:2016afs ; Millar:2018hkv , theoretical comparison to the dark matter-electron scattering Baxter:2019pnz ; Essig:2019xkx , the Migdal-photoabsorption relation Liu:2020pat , formulation for semiconductor response Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu , and the semi-inclusive calculation Cox:2022ekg .
@@ 2 Energy eigenstates of atomic system
where $\vec{p}_{n}$ and $\vec{p}_{e,i}$ is the three-dimensional momentum of the nucleus and $i$-th electron, $\vec{r}$ is the three-dimensional coordinates, $m_{n}$ and $m_{e}$ are the mass of the nucleus and electron, $Z$ is the atomic number, and $\alpha$ is the fine structure constant. The bracket $\{i,j\}$ represents the sum of the pair $i$ and $j$ under $i<j$.
@@
The above Hamiltonian can be separated by the center of mass energy and the others. Therefore, the energy eigenstates of the atom are given by $|\Psi_{A}\rangle=|\vec{p}_{\rm CM}\rangle\otimes|\Psi_{\rm re}\rangle$. Here, $|\Psi_{\rm re}\rangle$ is the energy eigenstate of the Hamiltonian for the relative coordinates
@@
and $|\vec{p}_{\rm CM}\rangle$ is the plane wave eigenstate for the center of mass state. Note that $H_{\rm re}$ includes the potential of nucleus being at the origin by approximation. Therefore, the relative coordinates represent the positions of electrons.
In the following, the energy eigenstate $|\Psi_{\rm re}\rangle$ is assumed to be given by the single Slater determinant of the single-electron wave function. Also, the single-electron wave functions are expressed by the spherical spinors (or spinor spherical harmonics) (see, e.g., Ref. johnson2007atomic for review).
@@ 3 Interaction potential
The interaction with the nucleus can be ignored approximately when the excitation and ionization processes are considered. This is because the potential from the nucleus interactions depend on the coordinates as $\vec{r}-\vec{r}_{n}$ and $\vec{r}_{n}$ is related to relative coordinates $\vec{r}_{i}$ with the suppression factor of $\approx m_{e}/m_{n}$.
@@
The above assumptions of a single Slater determinant and pure Coulomb potential neglect the electron correlation. These assumptions are not good in low energy scattering region and better in high energy region peixoto1969elastic ; naon1972importance ; khatun2021theoretical ; parvin2023theoretical . This is because the incident electron passes through the atom in a short time. Ref. peixoto1969elastic shows that those approximations look good for $40$ keV of incident energy and the discrepancy is about 6-12 % in small scattering angle region and better in large angle region for He and Ne atoms. In addition, this approach partially incorporates electron screening and correlation effects beyond the pure Coulomb potential and goes beyond a hydrogenic approximation.
@@ 4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in},s_{\rm in}\rangle_{I/F}$. Incoming electron states are taken as plane waves. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
@@
$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I},s_{{\rm in},I}\rangle$
$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F},s_{{\rm in},F}|\vec{r}\rangle e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}).$ (7)
@@
$\times\left(\frac{m_{A}}{m_{e}}\right)^{2}(2\pi)^{4}\delta^{4}(p_{{\rm in},F}+p_{{\rm CM},F}-p_{{\rm in},I}-p_{{\rm CM},I})$
$\times\left|\delta_{s_{\rm in}^{F}s_{\rm in}^{I}}\mathcal{M}_{t}(q^{2})Z_{FI}(\vec{q})+\delta_{s_{i}^{F}s_{\rm in}^{I}}\mathcal{M}_{u}(q_{u}^{2})Z_{{\rm in}I}(\vec{q}_{u})\right|^{2}.$ (12)
@@
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the difference of ionization cross section is around 30%30\ $\%$ from the database value in 5 keV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around 60%60\ $\%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the difference of ionization cross section is around 80 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This gap dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term is effective at low energy and negligible at high energy. Note that the probability is almost dominated by the excitation/ionization from the outermost shell. For Ar, $3p$ states dominate the contribution and $2p$ and $3s$ states give $\sim 1\ \%$ contribution at $E_{{\rm kin},I}=10$ keV. For Xe, $5p$ states dominate up to $\sim$ 100 eV, but, $4d$ states have $\sim 40\ \%$ contribution at $E_{{\rm kin},I}=1$ keV.
@@ 5 Comparison with other results
where $\beta$ is the velocity of incoming electron and the values of $M_{\rm Salvat}^{2}$ and $C_{\rm Salvat}$ are taken from the figure in Ref. salvat2022inelastic . This calculation used the shell corrections, self-consistent Dirac-Hartree-Fock-Slater potential, independent-electron approximation, and relativistic plane-wave Born approximation. They reported less than 1 % precision for $E\gtrsim 100$ MeV and $\sim 8$ % precision for $E\sim 5$ MeV for high $Z$ atoms and good agreement with the results of Ref. inokuti1975total ; inokuti1981oscillator .
@@
The binding energies and wave functions of electrons are calculated with electron correlations, e.g., Ref. Baxter:2019pnz shows $\sim 5$ % precision for binding energies which is better than Ref. Ibe:2017yqa . Also, Ref. Cox:2022ekg shows $\sim 20$ % differences of ionization probabilities for xenon 5s and argon 3p states with a 20 keV ionized electron. These uncertainties are significant for outer shell states and this comparison method actually relies on the precision of outermost shell states. Note that the uncertainty becomes much smaller for inner shell electrons. Density functional theory may provide better precision of binding energies, but, it does not provide the accurate wave functions.
@@ 6 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The differences are $\sim 30\ \%$ for Ar at 10 keV and $\sim 80\ \%$ for Xe compared with the database value shown in Fig. 1. For Xe, the database value is limited and the differences may be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. This approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected differences for excitation are not reliable. The differences in the low energy threshold come from the kinematical constraint, $E_{{\rm kin},I}>2\Delta E$. In addition, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ differences from the experimental data Ibe:2017yqa . In this calculation, the change of the electron mass in $m_{A}$ by scattering is neglected. This change will cause a difference of $m_{e}/m_{n}\sim 10^{-4}$.
@@
This treatment of the Migdal effect in Ref. Ibe:2017yqa does not include explicit couplings between different electronic channels due to non-adiabatic transitions. The dimensionless order parameter is assumed from the time-dependent perturbation theory as $\kappa_{ab}=|\vec{d}_{ab}\cdot\vec{v}_{n}|/E_{ab}$. Here, $\vec{d}_{ab}$ is the transition matrix of states $a$ and $b$, $\vec{v}_{n}$ is the velocity of nuclear recoil, and $E_{ab}$ is the energy difference. Assuming $|\vec{d}_{ab}\cdot\vec{v}_{n}|\sim a_{0}^{-1}|\vec{v}_{n}|$, the parameter becomes $\kappa_{ab}\sim(7\ {\rm keV}/E_{ab})\sqrt{E_{R}/m_{n}}$. Then, $\kappa_{ab}\ll 1$ will be satisfied at least for $E_{R}\lesssim 1$ keV or $E_{ab}\gtrsim 100$ eV. A fully consistent treatment of the Migdal effect including multi-channel non-adiabatic couplings and time-dependent correlation effects would require a more sophisticated quantum-chemistry or time-dependent multi-electron approach, which will be a future improvement.